A biotechnology firm is planning its investment strategy for future products and research labs. A poll found that 15​% of a random sample of 1024 adults approved of attempts to clone a human. Use this information to complete parts a through e. ​a) Find the margin of error for this poll if we want 99​% confidence in our estimate of the percent of adults who approve of cloning humans. MEequals nothing ​(Round to three decimal places as​needed.) Expert Answer Anonymous Anonymous answered thisWas this answer helpful? 0 0 4,763 answers \hat{p} = 0.15 Sample size = n = 1024 Here ME is a margin of error. ME = ZC* (1-P) Zc = 2.58 ( From z table) ME = 2.58 * V 0.15*(1 – 0.15) - 258 1024 0.000125 = 0.029 Margin of error = ME = 0.029

​b) Explain what that margin of error means. Choose the correct answer below.

A.

The margin of error is the width of the confidence interval that contains the true proportion of adults who approve of attempts to clone a human.

B.The margin of error is the value that should be subtracted from the

9999​%

confidence level to obtain the​ pollsters' true confidence level.

C.The pollsters are

9999​%

confident that the true proportion of adults who approve of attempts to clone a human is within the margin of error of the estimated

1515​%.

D.The pollsters are

9999​%

confident that the margin of error contains the true proportion of adults who approve of attempts to clone a human.
Answer #1

clas Date Page P=0.15 n=1024 @ Given that =0.01 Margin of enor: Zay, JPCI-P3 = (2-545) 0.15X0495 = 12.575) (010112) = 0.029 A